Sue leaves home at 9 a.m. bicycling at a rate of 28 mi/h. Two hours later, Joe leaves driving at the rate of 48 mi/h. At what time will Joe catch up with Sue?

Answer by Paul(988)

Let t be the time
28(t+2)=48t
28t+56=48t
20t=56
t=2.8
9+2.8=11.8
Hence, Joe will catch sue at 11.8 am 2.8 hours from where sue left